OpenStudy (anonymous):

Find the amplitude of y = -4 sin x.

jimthompson5910 (jim_thompson5910):

In general y = A*sin(Bx-C) + D has an amplitude of |A|

jimthompson5910 (jim_thompson5910):

y = A*sin(Bx-C) + D |A| is the amplitude T = 2pi/B is the period C/B is the phase shift y = D is the midline

OpenStudy (anonymous):

so the answer is 4?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

okay thank you!